# JEL: C81, O32

### v3

@@ 3. Disentangling Trends in Publication Data
Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.282828Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial. We find similar patterns in PubMed as a whole, and in alternative cuts of records. Across fields, the total number of scientific papers, the total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals exhibit similar patterns. These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024). In this section, we use our census of clinical trials to decompose this trend. We contrast trends in the production of clinical trials with those documented for other, similar forms of medical research.
@@ 3.1. Decomposing Trends
These trends are robust to alternative forms of data construction. In Figure D.1, we construct these plots using our liberal and moderate sample definitions and find, qualitatively, no change. In Figure D.2, we explore a potential concern with a citation-based measure. We document in Figure 2.1 that the quantity of scientific publications is rising over time. It is reasonable to expect that this rise, coupled with the ease of citing papers enabled by the internet, coincides with an increase in the total number of citations in the aggregate and per paper. We confirm, empirically, that both hypotheses are borne out by our data. To account for the fact that the “cost” of citing a clinical trial may have decreased over time, thus generating the trend documented in Figure 3.1, we construct a version of this plot in which each citing paper is weighted by the total number of papers that it cites. That is, a citing paper that cites a trial and one other paper receives a weight of 1/2, and a citing paper that cites a trial and nine other papers receives a weight of 1/10. In doing so, we “penalize” papers that are especially generous in their tendency to cite other papers. Figure D.2 suggests, again, that this trend is robust to this alternate construction.
@@ 3.2. Clinical Trials as a Productivity Indicator
We view the first potential explanation as unlikely for two reasons. Somewhat suggestively, NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable over this period. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this measurement challenge. More directly, recent work by Sertkaya et al., (2024) suggests that the costs of drug development—measured using the contracts associated with particular study sites in clinical trials—have been essentially constant from 2000 to 2018. The second potential explanation is unlikely from a more conceptual perspective. We find essentially complete stability in this measure over time: any changes in publication bias or trial failure rates must, then, perfectly offset changes in public funding patterns, to yield the flat series here.
Figure 3.3 provides a more granular view of this heterogeneity. We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.323232Appendix D presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. Observe that we elongate the y-axis to highlight the right-tail of the citation distribution. The quantiles corresponding to the right tail of the citation distribution are constant across years. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
@@
Figure D.4 is constructed analogously, where instead we display quantiles of the distribution of the number of citations received from all publications indexed in PubMed. There, we find that the right-tail of this distribution is increasing over time. In other words, as the total number of publications is increasing, the total number of citations received by clinical trials is increasing. However, these citations are not originating in leading journals.
@@
We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period. There are four potential concerns with this interpretation.
@@
First, if the costs of either upstream, basic inputs into clinical trials or the costs of clinical trials themselves have increased substantially over this time, the total resources devoted toward research will have increased, even though Figure 3.1 suggests that counts of trials produced are stable. In our view, this is unlikely to be the case. On the issue of clinical trial costs, the most up-to-date, comprehensive evidence on drug development costs, in our read, is given in Sertkaya et al., (2024), who use internal data from the U.S. Food and Drug Administration and contracts associated with clinical trial sites to compute estimates for the period from 2000 to 2018. In these data, they find that over this period, “the cost of drug development remained relatively stable or may have even decreased.” [^33] This finding revises a frequently cited claim made in popular press interviews (e.g., Herper,, 2022), in academic articles that rely on proprietary data sources (e.g., DiMasi et al.,, 2016), and in surveys on research productivity (Goldin et al.,, 2024)—which suggest that clinical trial and drug development costs are sharply increasing over time. This revision is in the spirit of an argument made by Cockburn, (2006), in response to an earlier debate about drug development costs. Thus, we do not consider the prospect of rising study prices a concern for the interpretation of this pattern. On the issue of basic science inputs into clinical trials, there is little empirical evidence that increases in cost or decreases in the productivity have occurred. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science. In fact, changes in the production of basic scientific inputs around the year 2000—enabled by breakthroughs such as the Human Genome Project and the advent of high-throughput screening for drug discovery (see Scannell et al.,, 2012)—suggest, if anything, that costs of basic science inputs have fallen. Moreover, it is reasonable that such trends would affect trends in trial quantity, quality, composition, or cost. Given the absence of such patterns, it is unlikely that changes in basic science threaten the interpretation of our findings. [^34] Constructing a comprehensive measure of medical research productivity, which grapples with these various input markets, remains an open question. To our knowledge, the work that is closest in spirit is Berndt and Cockburn, (2013), who aim to develop a price index for clinical trial research, with the eventual goal of developing price indices for biomedical research more generally.
@@
Second, one might be concerned about how to interpret the x-axis in our plots. In particular, the relationship between clinical trial completion dates and clinical trial start dates varies systematically within and across disease classes (Budish et al.,, 2015). Here, we follow existing literatures that use dates affixed to patents and publications as measures of the timing of research. Practically speaking, as Footnote 14 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, are available only with a considerable lag. We focus on the date of publication, then, for ease of interpretation.
@@
Third, our findings are threatened by publication bias. The existence of publication bias means that counts of scientific publications only partially measure the quantity of research (Andrews and Kasy,, 2019). Our main finding is that the quantity of clinical trials is stable over time. If a trend were in fact present, and our measure were confounded by publication bias, then it would need to be the case that the trend was perfectly offset by the bias. It is reasonable to view cancellation, in this way, as unlikely.
@@
Fourth, and finally, we must assume that the number of publications associated with each clinical trial is constant across our period of interest. We view this assumption as reasonable for three reasons. First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010). Under these rules, there is little scope for authors to, for example, split results across several publications. Second, when constructing our sample of clinical trials, we have excluded publications that re-analyze data generated by an existing study. Third, more suggestively, we find no evidence of a change in this relationship between the size of NIH grants and the quantity of resulting publications reporting the results of a clinical trial during our period of interest. [^35] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019. Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials. For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial. This is stable across years of data.
@@ 3.3. Why is the Quantity of Research Rising?
The quantity of papers citing clinical trials has increased, per Figure 3.1, by nearly a factor of two over our period of interest. In this section, we examine various measures of publication composition and quality to shed light on the source of this increase. Our aim is to determine the source of the categorization errors that yield measures such as those in Figure 2.1. In doing so, we provide suggestive evidence on whether measured increases in the quantity of scientific research correspond to “real” increases in the quantity of underlying research or whether they capture changes in the ease, frequency, and composition of publication.
@@ 3.3.3. Publication Quality
Next, we collect a set of facts about the quality of this growing set of “citing” papers. We begin in Figure 3.5 by replicating Figure 3.3 for the sample of papers that cite clinical trials. There are several elements of this plot that are informative. First, the share of “citing” papers with a small number of citations appears to grow over time. That is—the left tail of the citation distribution for this set is both long and growing. Note also that the right tail of the citation distribution appears to shrink over time. That is, the proportion of highly cited papers appears to be shrinking. These patterns provide one indication that the central tendency of the quality of non-clinical-trial medical research may be decreasing over time. [^44] Figure D.5 gives an analogous figure, displaying the distribution in citations from all publications to the sample of publications that cite clinical trials. Here, the right tail of the distribution of citations is growing.
@@ 4. Discussion
Our findings focus on one aspect of the larger biomedical research ecosystem: investments in efforts to bring basic science discoveries to market, in the form of new medicines. Development of supplemental measures of medical research productivity, which capture the efficiency of basic scientific research, is a useful area for future research.
@@
Supplemental Appendix to: Distilling Data from Large Language Models:
An Application to Research Productivity Measurement [^1] Date: September 11, 2024
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
@@ Appendix A Publication Data
In this Appendix, we detail our treatment of the PubMed / MEDLINE database (hereafter, “PubMed”). We give a general overview of the PubMed data in Section A.1. In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 3.1. We discuss the process of determining the set of leading medical journals in Section A.3.
@@ A.1. The PubMed / MEDLINE Database
Our version of the PubMed / MEDLINE database contains roughly $34$ million records and is current through December 2022. These records were constructed by parsing bulk MEDLINE XML files. [^48] We are grateful to Heidi Williams for sharing this processed data. Technically, PubMed and MEDLINE are different products. MEDLINE, a subset of PubMed, is the U.S. National Library of Medicine’s (NLM) bibliographic database, which contains references to journal articles in life sciences, with a primary focus on biomedicine. A committee at the NLM determines the set of indexed journals, meaning that only journals that meet certain quality and content standards are indexed. In practice, this means that so-called predatory journals are excluded, as are pre-prints and non-peer reviewed articles. Informal conversations with staff at the NLM suggest that MEDLINE should contain the universe of peer-reviewed publications in legitimate journals. PubMed includes a broader set of records, including pre-prints and publications deposited through alternative processes. Official documentation for MEDLINE is available at the link: https://www.nlm.nih.gov/databases/download/pubmed_medline_documentation.html. Throughout the text, and from this point forward in the Appendix, we refer to this database as “PubMed.”
A natural question for researchers who rely on these data as a census of research investments is whether PubMed records are complete and accurate. To our knowledge, there is no paper that reports such validation exercises for each field of PubMed. We focus on those data elements most relevant to our work. There are 2,335,653 records in PubMed that have no associated year of publication, which we drop from our baseline sample. There are 16 records, across all years of data, that are not indexed with tags describing their contents (NML tags). From 2010 forward, roughly 92 percent of records have associated abstracts. We randomly inspect roughly 150 records with no abstracts. None correspond to publications that satisfy our definition of a clinical trial.
We assume, throughout this paper, that records with missing citation data have zero associated citations. We confirm that this is generally accurate in two ways. First, we search for Google Scholar records corresponding to roughly 100 randomly selected records with zero citations. In each case, Google Scholar indicates that the paper has no more than two citations. Second, we link records in PubMed to Clarivate Analytics’ Web of Science database. We compare citation counts constructed using information in the two databases. We find that the two measures have a correlation of 0.9.
Researchers interested in different cuts of the data may need to conduct additional validation exercises. Figure 1 highlights that the frequencies of certain NLM tags have changed over time, often sharply. Researchers should account for such changes in any use of these flags. For other assessments of the completeness of certain PubMed fields, see Durvasula et al., (2021) and Ouellette and Sampat, (2024).

### v5

@@ 3. Disentangling Trends in Publication Data
Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.363636Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial. We find similar patterns in PubMed as a whole, and in alternative cuts of records. Across fields, the total number of scientific papers, the total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals exhibit similar patterns. These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024). In this section, we use our census of clinical trials to decompose this trend. We contrast trends in the production of clinical trials with those documented for other, similar forms of medical research.
@@ 3.2. Clinical Trials as a Productivity Indicator
We view the first potential explanation as unlikely for two reasons. Somewhat suggestively, NIH RePORTER data allow us to tabulate total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable over this period. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this measurement challenge. More directly, recent work by Sertkaya et al., (2024) suggests that the costs of drug development—measured using the contracts associated with particular study sites in clinical trials—have been essentially constant from 2000 to 2018. The second potential explanation is unlikely from a more conceptual perspective. We find essentially complete stability in this measure over time: any changes in publication bias or trial failure rates must, then, perfectly offset changes in public funding patterns, to yield the flat series here.
Figure 3.3 provides a more granular view of this heterogeneity. We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.404040Appendix D presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. Observe that we elongate the y-axis to highlight the right-tail of the citation distribution. The quantiles corresponding to the right tail of the citation distribution are constant across years. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
@@
We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period. At the very least, we find that there is insufficient evidence to reject the null that pharmaceutical research productivity has been stable over this time period. There are four potential concerns with this interpretation.
@@
First, if the costs of clinical trials have increased substantially during this time, the total resources devoted to the commercialization of drugs will have increased, although Figure 3.1 suggests that the quantity of trials produced has been stable. In our view, this is unlikely to be the case. The most up-to-date, comprehensive evidence on clinical trial and drug development costs is given in Sertkaya et al., (2024), which uses internal data from the U.S. Food and Drug Administration and contracts associated with clinical trial sites to compute estimates for the period from 2000 to 2018. In these data, they find that over this period, “the cost of drug development remained relatively stable or may have even decreased.” [^41] This finding revises a frequently cited claim made in popular press interviews (e.g., Herper,, 2022), in academic articles that rely on proprietary data sources (e.g., DiMasi et al.,, 2016), and in surveys on research productivity (Goldin et al.,, 2024)—which suggest that clinical trial and drug development costs are sharply increasing over time. This revision is in the spirit of an argument made by Cockburn, (2006), in response to an earlier debate about drug development costs. Thus, we do not consider the prospect of rising study prices a concern for the interpretation of this pattern.
A distinct and somewhat related concern is that the cost of basic scientific research “upstream” of clinical trials could have increased substantially, either because more studies are necessary to produce a candidate drug or because the cost of inputs has increased. To be clear, our interest in this paper is on the productivity of the pharmaceutical industry, not biomedical research as a whole. Debates about the productivity of the pharmaceutical industry hinge on the idea that efforts to commercialize medicines are slowing—because firms are too inefficiently managed to run new studies (Ruffolo,, 2006), because concerns about commercial viability encourage firms to shelve valuable drug candidates (Scannell et al.,, 2012), or because regulation has become too onerous (Scannell et al.,, 2012). Changes in the productivity of basic science are interesting and important, to be sure, but not the central object.
Nonetheless, in our read, there is essentially no empirical evidence indicating either decreases in the productivity of basic scientific studies or increases in their cost. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in the quantity of inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science. In fact, changes in the production of basic scientific inputs around the year 2000—enabled by breakthroughs such as the Human Genome Project and the advent of high-throughput screening for drug discovery (see Scannell et al.,, 2012)—suggest, if anything, that costs of basic science inputs have fallen. In fact, the pharmaceutical productivity literature often characterizes the productivity “puzzle” in terms of this decrease in costs (see e.g., Cockburn,, 2006).
There is some evidence that the total quantity of scientific papers has increased over this time period, which has been interpreted in turn as evidence of increasing (basic) scientific research effort—notably in Bloom et al., (2020) and Park et al., (2023). In our view, moving from these trends in publication data to a conclusion of decreasing productivity is premature. In the next section, we document that measured increases in the quantity of scientific publications capture, in substantial part, a sharp increase in the quantity of low-quality papers over time. [^42] Park et al., (2023) document the same: in their data, the quantity of low-impact research has increased substantially, while the quantity of high-impact research has remained essentially stable over time. Their measure of impact is a measure of “disruptiveness”—how much a given piece of research deviates from existing work. If measured increases in the quantity of papers reflect, say, an increase in the likelihood that the marginal paper is published, perhaps because the cost of publication has decreased, publication trends alone are largely uninformative on the question of scientific productivity.
Our data also provide a more suggestive piece of evidence on trends in upstream research. We find remarkable stability in trial quantity, quality, and composition. Sertkaya et al., (2024) finds analogous stability in cost. If there were large changes in upstream productivity or cost, we might expect to see some indication in these downstream markets. In the absence of such indicators, we conclude that large shifts in the productivity of upstream markets are unlikely to be relevant considerations for the interpretation of our findings. [^43] Constructing a comprehensive measure of biomedical research productivity, which grapples with these various input markets, remains an open, important challenge. Our work does not speak directly to this challenge, but does highlight limitations in existing measures of productivity and challenges with the use of publication data.
@@
Second, one might be concerned about how to interpret the x-axis in our plots. In particular, the relationship between clinical trial completion dates and clinical trial start dates varies systematically within and across disease classes (Budish et al.,, 2015). Here, we follow existing papers that use completion dates affixed to patents and publications as measures of the timing of research. In the context of clinical trials, the use of completion dates is reasonable as a practical matter. As Footnote 21 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, available only with a considerable lag. The date of publication is more easily interpretable.
Third, our findings may be threatened by publication bias. Specifically, the existence of publication bias means that counts of scientific publications only partially measure the quantity of clinical trials being produced in each time period (Andrews and Kasy,, 2019). To weigh concerns about publication bias, it is useful to recall that our objective is to produce an estimate of trends in clinical trial production, sufficient to test existing hypotheses about declining productivity in this industry. Our aim is not to produce a comprehensive census of all clinical trials. Given this objective, publication bias is a serious threat to the interpretation of our results if it is meaningfully changing over this time period.
Recent empirical evidence in Oostrom, (2024) suggests that publication bias may have been stable over this time period. Oostrom collects a set of roughly 600 clinical trials associated with psychiatric medications, as listed in a comprehensive meta-analysis spanning five decades. For each record, the author collects any associated scientific publications and any ClinicalTrials.gov entries. Of interest to our analysis are two facts in this case study. First, Oostrom documents evidence of publication bias across years of her data, but finds that trends in this bias become stable by the early 2000s. [^44] See Oostrom, (2024), Figure5(b). That is, although there is publication bias, it is not meaningfully changing in her sample of trials across our period of interest. Second, she finds that trends in ClinicalTrials.gov reporting shift sharply over our period of interest. In 2010, roughly 40 percent of the trials in her sample were registered on ClinicalTrials.gov. By 2019, roughly 80 percent were registered. Taken together, her findings lend credence to our assumption that publication bias and publication reporting norms are relatively stable during this period, whereas database reporting is not.
@@
In our data, we find that the quantity, quality, and composition of published clinical trials have been stable over time. If a trend were in fact present, and our measure were confounded by publication bias, then it would need to be the case that the trend was perfectly offset by the bias. Cancellation in this way—across all three margins we consider—is unlikely.
@@
Fourth, and finally, we must assume that the number of publications associated with each clinical trial is constant across our period of interest. We view this assumption as reasonable for three reasons. First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010). Under these rules, there is little scope for authors to, for example, split results across several publications. Second, when constructing our sample of clinical trials, we excluded publications that re-analyze data generated by an existing study. Third, more suggestively, we find no evidence of a change in this relationship between the size of NIH grants and the quantity of resulting publications reporting the results of a clinical trial during our period of interest. [^45] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019. Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials. For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial. This is stable across years of data.
@@ 3.3. Why is the Quantity of Research Rising?
The quantity of papers citing clinical trials has increased, per Figure 3.1, by nearly a factor of two across our period of interest. In this section, we examine various measures of publication composition and quality to shed light on the source of this increase. Our aim is to determine the source of the categorization errors that yield measures such as those in Figure 2.1. In doing so, we provide suggestive evidence on whether measured increases in the quantity of scientific research correspond to “real” increases in the quantity of underlying research or whether they capture changes in the ease, frequency, and composition of publication.
@@ 3.3.3. Publication Quality
Next, we collect a set of facts about the quality of this growing set of “citing” papers. We begin in Figure 3.5 by replicating Figure 3.3 for the sample of papers that cite clinical trials. There are several elements of this plot that are informative. First, the share of “citing” papers with a small number of citations appears to grow over time. That is—the left tail of the citation distribution for this set is both long and growing. Note also that the right tail of the citation distribution appears to shrink over time. That is, the proportion of highly cited papers appears to be shrinking. These patterns provide one indication that the central tendency of the quality of non-clinical-trial medical research may be decreasing over time.
@@ 4. Discussion
Our findings focus on one aspect of the larger biomedical research ecosystem: investments in commercialization efforts in the pharmaceutical industry, which bring products to market in the form of new medicines. Development of new measures of upstream research productivity—centered on basic discoveries in biomedical research and translational efforts to map these discoveries to products safe for testing in human subjects—are a natural area for future research.
@@
Supplemental Appendix to: Counting Clinical Trials:
New Evidence on Pharmaceutical Sector Productivity [^1] Date: January 24, 2025
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3appendix.Dsubsection.D.1subsection.D.2subsubsection.D.2.1subsubsection.D.2.2subsubsection.D.2.3appendix.E
@@ A.1. The PubMed / MEDLINE Database
Our version of the PubMed / MEDLINE database contains roughly $34$ million records and is current through December 2022. These records were constructed by parsing bulk MEDLINE XML files. [^57] We are grateful to Heidi Williams for sharing this processed data. Technically, PubMed and MEDLINE are different products. MEDLINE, a subset of PubMed, is the U.S. National Library of Medicine’s (NLM) bibliographic database, which contains references to journal articles in life sciences, with a primary focus on biomedicine. A committee at the NLM determines the set of indexed journals, meaning that only journals that meet certain quality and content standards are indexed. In practice, this means that so-called predatory journals are excluded, as are pre-prints and non-peer reviewed articles. PubMed includes a broader set of records, including pre-prints and publications deposited through alternative processes. Official documentation for MEDLINE is available at the link: https://www.nlm.nih.gov/databases/download/pubmed_medline_documentation.html. Throughout the text, and from this point forward in the Appendix, we refer to this database as “PubMed.”
There are 16 records, across all years of data, that are not indexed with tags describing their contents (NML tags). From 2010 forward, roughly 92 percent of records have associated abstracts. We randomly inspect roughly 150 records with no abstracts. None correspond to publications that satisfy our definition of a clinical trial. We assume, throughout this paper, that records with missing citation data have zero associated citations. We confirm that this is generally accurate in two ways. First, we search for Google Scholar records corresponding to roughly 100 randomly selected records with zero citations. In each case, Google Scholar indicates that the paper has no more than two citations. Second, we link records in PubMed to Clarivate Analytics’ Web of Science database. We compare citation counts constructed using information in the two databases. We find that the two measures have a correlation of 0.9.